# MARBLE-Net: Learning to Localize in Multipath Environment with Adaptive Rainbow Beams

### v1

@@ I Introduction
The evolution toward 6G networks features the deep integration of sensing and communication, giving rise to the paradigm of Integrated Sensing and Communication (ISAC) [Gonzalez-Prelcic2025Six]. Leveraging large bandwidths and massive antenna arrays, ISAC enables high-precision localization as a native network service [Jabeen2025Localization]. However, this integration exposes a dichotomy in how wireless systems interact with multipath propagation. For communications, multipath is a valuable resource that enables diversity and multiplexing gains [Alkhateeb2014MIMO]; for monostatic sensing, it is often regarded as interference that causes ambiguities and degrades performance [jia2024multipath, lu2022degrees].
@@
This challenge has spurred a paradigm shift from multipath mitigation to multipath exploitation, where multipath components are reinterpreted as rich sources of information. In multipath-rich environments, echoes via non-line-of-sight (NLoS) paths provide additional observation perspectives and spatial diversity, improving sensing accuracy, particularly for targets with fluctuating radar cross-sections (RCS) [zhang2024multipath, lv2025target]. For extended targets such as vehicles, multipath echoes can even exceed direct reflections and carry more precise motion information [zhao2025sensing]. The goal is thus to transform multipath “from enemy into friend” by exploiting its embedded geometric cues for localization tasks [liu2025multipath].
@@
A highly efficient physical-layer mechanism for spatial exploration in wideband ISAC systems is frequency-dependent beamforming, which forms “rainbow beams” through controlled beam squint[Yan2019Wideband, luo2023beam, Cui2023NearField, gao2023integrated, Zhou_Gui_JSAC25, Zheng2025Near]. By employing true-time-delay (TTD) devices, often as part of a phase-time array (PTA) [ratnam2022joint, Mo_Jianhua_Asilomar24, cai2025hybrid, Liang2025CFARNet, nam2025joint], this technique uniquely maps different frequency subcarriers to distinct spatial locations, enabling a wide angular sector to be scanned within a single OFDM symbol. Such “single-scan” operation greatly reduces sensing overhead and latency, making it ideal for real-time applications. The beam squint effect, usually viewed as a communication impairment, is thus harnessed to encode angular information directly in the frequency domain.
@@
Despite this efficiency, interpreting the complex multipath-rich echoes from rainbow beams remains challenging [mokri2024custom]. The superposition of numerous propagation paths forms a unique frequency-domain “fingerprint” for each target, too intricate for conventional estimators that rely on dominant LoS paths. This complexity, however, creates an opportunity for data-driven methods. Deep learning (DL) excels at modeling non-linear mappings from high-dimensional signals to geometric outputs [papageorgiou2021deep, chen2024sdoa]. While prior works apply DL for DoA estimation [chen2024sdoa] or near-field localization with fixed rainbow beams [klus2025deep], jointly leveraging multipath propagation and adaptive beam squint within a unified neural framework remains largely unexplored.
@@
Most existing methods, including DL-based ones, rely on fixed, pre-designed rainbow beams that are non-adaptive and often suboptimal for specific environments [klus2025deep]. To overcome this limitation, we adopt a co-design paradigm that jointly optimizes the sensing beam and estimation network. By treating the beamforming parameters—phase-shifter (PS) and TTD values—as learnable weights in an end-to-end framework, the system autonomously shapes its rainbow beams and exploits environmental multipath information. Building on this idea, we propose a novel deep learning framework for single-target localization that jointly optimizes the parameters of a TTD-based rainbow beamformer and a subsequent convolutional neural network (CNN) for position estimation. Our key contributions are summarized as follows:
@@
  • Synergistic Co-design for Adaptive Sensing: We propose a fully differentiable, end-to-end model, named MARBLE-Net (Multipath-Aware Rainbow Beam LEarning Network), that jointly optimizes a learnable rainbow beamformer and a localization CNN. This co-design enables the system to adapt its physical-layer sensing beam to the specific propagation environment via gradient-based optimization.
  • Structured Training Methodology: We propose and validate a multi-stage training strategy, which first pre-trains the prediction network with a fixed beamformer, then adapts the beamformer’s physical parameters, and concludes with an end-to-end fine-tuning. This structured approach is critical for stable convergence and superior performance. Extensive simulations show that MARBLE-Net achieves lower localization error compared with fixed-beam deep learning and traditional k-NN methods.
@@ II-A Rainbow Beamforming
$\mathbf{w}_{m}]_{n}=\frac{1}{\sqrt{N_{t}}}\exp\left(-\mathrm{j}\left(\phi_{n}+2\pi(f_{m}-f_{0})\tau_{n}\right)\right),$ (1)
where $f_{m}$ is the frequency of the $m$-th subcarrier and $f_{0}$ is the carrier frequency, $\phi_{n}$ and $\tau_{n}$ are the phase shift and time delay applied at the $n$-th antenna. In our framework, the phase shift and time delay values are learnable parameters. They are initialized based on the strategy in [luo2023beam] to span a predefined angular range but are subsequently optimized during training.
@@
where $\mathbf{h}_{m}\in\mathbb{C}^{N_{r}\times 1}$ is the channel vector, $P_{t}$ is the transmit power of UE, $x_{m}$ is the transmitted symbol, and $\eta_{m}$ is i.i.d. complex additive white Gaussian noise (AWGN). The collection of received samples, $\mathbf{y}=[y_{1},\dots,y_{M}]^{T}\in\mathbb{C}^{M\times 1}$, forms the observation vector.
@@ II-B High-Fidelity Ray-Tracing Channel Generation
    LoS-only Scene: An unobstructed environment with no obstacles, where propagation is dominated by the direct LoS path. This serves as a baseline reference.
@@
    ‘circle’ Scene: An environment designed to generate uni-directional multipath with low angular spread. As shown in Fig. 2(a), it consists of a semicircular wall (radius $300\text{\,}\mathrm{m}$, height $50\text{\,}\mathrm{m}$), which confines most reflected paths to a narrow azimuth range. Although these paths may differ in elevation due to wall height and target position, they predominantly arrive from the same horizontal direction, forming a characteristic uni-directional multipath pattern.
@@
    ‘rounded-L’ Scene: An environment designed to generate multipath with high angular spread. As shown in Fig. 2(b), it consists of a $300\text{\,}\mathrm{m}$ radius quarter-circle arc contiguous with a $400\text{\,}\mathrm{m}$ straight wall (both $50\text{\,}\mathrm{m}$ high). This structure introduces reflective paths from more diverse directions compared to the ‘circle’ scene.
@@
    ‘L’ scene: An environment creating a rich multipath profile including higher-order reflections. It is composed of two perpendicular walls (North-South wall at $205\text{\,}\mathrm{m}$ from origin, East-West wall at $180\text{\,}\mathrm{m}$ from origin, both $50\text{\,}\mathrm{m}$ high). This geometry is representative of urban street-corner scenarios, generating both single- and double-bounce reflection paths.
For all scenes, the radio material properties are assigned to be concrete for the ground and brick for the walls. The Blender scene is imported into Sionna, and for each of the 100 000100\,000 random UAV positions, Sionna’s path solver traces interactions (reflections, refractions) to compute all valid propagation paths and synthesize the final frequency-domain channel matrix $\mathbf{H}$.
@@ III-A Learnable Rainbow Beamformer ($\mathcal{M}_{1}$)
This module acts as the physical-layer front end. Its trainable parameters consist solely of the PS and TTD vectors, $\bm{\phi},\bm{\tau}\in\mathbb{R}^{N_{r}}$. Given the channel matrix, it computes the received spectrum $\mathbf{y}$ using (1) and (2), and produces the log-magnitude power spectrum $\mathbf{P}_{\text{dB}}\in\mathbb{R}^{M}$ with each element defined as $\mathbf{P}_{\text{dB},m}=10\log_{10}(|y_{m}|^{2}+\epsilon)$. This logarithmic transformation effectively compresses the wide dynamic range of wireless signals, facilitating stable learning.
@@ III-B Prediction Network ($\mathcal{M}_{2}$)
The prediction network is a 1D CNN designed to regress the target’s 2D Cartesian coordinates from the 1D power spectrum.
@@ III-C Multi-Stage Training Strategy
Optimizing a coupled system of a physical-layer model and a neural network presents unique challenges due to the complex, non-convex loss landscape. To ensure stable convergence and achieve superior performance, we propose a structured multi-stage training strategy, akin to curriculum learning. This approach methodically optimizes different parts of the MARBLE-Net.
  • Stage 1: Prediction Network Pre-training. Train only the prediction network ($\mathcal{M}_{2}$) while keeping the beamformer ($\mathcal{M}_{1}$) fixed with its initial parameters. This phase allows the network to learn a mapping from a non-adaptive, but structured, input spectrum to the target’s location, establishing a strong feature extraction baseline.
  • Stage 2: Beamformer Adaptation. Freeze the weights of the prediction network ($\mathcal{M}_{2}$) and train only the learnable PS and TTD parameters of the beamformer ($\mathcal{M}_{1}$). In this stage, the beamformer is forced to adapt its physical sensing pattern to generate a power spectrum that is more “interpretable” for the pre-trained network, thereby minimizing the downstream localization error.
  • Stage 3: End-to-End Fine-tuning. Unfreeze all parameters and jointly train the entire MARBLE-Net model. This allows for fine-grained adjustments between the beamformer and the prediction network, leading to a synergistic optimization of the end-to-end system.
@@ IV-A Simulation Setup
Scene Metric k-NN RaiNet [klus2025deep] RaiNet MARBLE-Net MARBLE-Net
@@
LoS-only Localization RMSE (m) 15.73 4.02 3.62 3.94 3.46
Angle RMSE () 0.13 1.21 1.11 1.55 1.17
Range RMSE (m) 15.73 3.82 3.44 3.84 3.16
‘circle’ Localization RMSE (m) 16.63 5.34 4.94 4.77 4.28
Angle RMSE () 0.18 1.61 1.47 1.36 5.49
Range RMSE (m) 16.63 5.15 4.78 4.65 4.06
‘rounded-L’ Localization RMSE (m) 13.32 4.18 3.75 3.93 2.96
Angle RMSE () 0.17 1.25 1.20 1.51 1.41
Range RMSE (m) 13.32 3.99 3.51 3.71 2.78
‘L’ Localization RMSE (m) 5.85 1.88 1.55 1.06 0.88
Angle RMSE () 0.14 0.71 0.54 0.43 0.35
Range RMSE (m) 5.84 1.63 1.38 0.94 0.81
@@ IV-B2 RaiNet
As a point of comparison, the baseline RaiNet given in [klus2025deep] adopted a fixed rainbow beam and only trained the prediction network. This represents a conventional DL-based approach with a non-adaptive sensing beam. The RaiNet architecture consists of a feature extractor and a classifier. The feature extractor contains three 1D convolutional layers, followed by a classifier with three fully connected layers. The Tanh function is used as the primary activation function throughout the network.
@@ IV-C Performance Comparison and Ablation Study
First, it is evident that all deep learning approaches dramatically outperform the traditional k-NN baseline. In the ‘circle’ scene, for instance, k-NN yields a large localization RMSE of $16.63\text{\,}\mathrm{m}$, underscoring its inability to handle the complex signal fingerprints created by multipath propagation. Template-matching methods like k-NN fail when the direct relationship between the signal spectrum and geometric location is severely distorted, rendering the pre-compiled codebook ineffective.
@@
To deconstruct the sources of performance gain, we compare four DL-based configurations: RaiNet with a fixed beamformer (‘RaiNet (fixed)’), RaiNet with multi-stage training (‘RaiNet (adaptive)’), MARBLE-Net’s architecture with a fixed beamformer (‘MARBLE-Net (fixed)’), and the fully optimized MARBLE-Net. This allows us to isolate the effects of network architecture and joint training.
The Benefit of Joint Training: A key finding is the critical importance of joint end-to-end training. Comparing the ‘fixed’ and ‘multi-stage’ trained results for both architectures reveals a consistent performance improvement. For RaiNet in the ‘circle’ scene, multi-stage training reduces the RMSE from $5.34\text{\,}\mathrm{m}$ to $4.94\text{\,}\mathrm{m}$, a relative improvement of 7.5%. However, the benefit is far more pronounced for MARBLE-Net, where the RMSE drops from $4.77\text{\,}\mathrm{m}$ to $4.28\text{\,}\mathrm{m}$, a substantial 10.3% improvement. This demonstrates that while a simple network like RaiNet gains moderately from adapting the beamformer, the more complex architecture of MARBLE-Net synergizes much more effectively with the adaptive physical layer, unlocking noticeable performance gains.
@@
The Nuanced Impact of Multipath: The results in Table II reveal a complex, non-monotonic relationship between multipath complexity and localization accuracy. When moving from the ‘LoS-only’ scene (MARBLE-Net RMSE of $3.46\text{\,}\mathrm{m}$) to the highly complex ‘circle’ scene, performance for all models degrades. For MARBLE-Net, the RMSE increases to $4.28\text{\,}\mathrm{m}$. This suggests that the dense, uni-directional multipath in the ‘circle’ scene introduces ambiguity that is difficult to resolve, even for an adaptive system.
However, the ‘rounded-L’ scene tells a different story. Counter-intuitively, MARBLE-Net’s performance in this moderately complex scene (RMSE of $2.96\text{\,}\mathrm{m}$) is superior to both the ‘LoS-only’ and ‘circle’ scenes. This strongly indicates that MARBLE-Net is not merely mitigating multipath but is actively exploiting the structured, directional NLoS paths provided by the ‘rounded-L’ scene’s geometry. These paths act as a source of valuable spatial diversity, allowing the model to refine its position estimate. In contrast, the fixed-beam models fail to leverage this, showing worse performance in the ‘rounded-L’ scene than in the LoS scene, confirming that adaptation is the key to turning specific multipath from an impairment into a resource.
This trend of multipath exploitation is further exemplified by the ‘L’ scene. In this environment with two perpendicular walls, the localization error for MARBLE-Net drops to $0.88\text{\,}\mathrm{m}$. This is a significant improvement over all other scenarios, including LoS-only ($3.46\text{\,}\mathrm{m}$). This result suggests that the simple, planar reflections from the ‘L’ scene provide highly stable and geometrically informative NLoS paths, which the adaptive beamformer and network learn to use as ‘virtual anchors’, dramatically refining the position estimate.
The Benefit of Network Complexity: Finally, the role of network architecture is highlighted. With its fixed beamformer, MARBLE-Net’s more complex architecture (‘MARBLE-Net (fixed)’) offers a consistent but marginal advantage over the simpler ‘RaiNet (fixed)’ across all scenes. This suggests that simply increasing the number of network parameters is insufficient. It is the combination of a high-capacity prediction network and a co-adapted, learnable physical-layer beamformer that enables the system to learn the intricate mapping from multipath-rich channel observations to precise location estimates, unlocking the large performance leap seen in the ‘MARBLE-Net (adaptive)’ results.
Figure 5: CDF of localization error for MARBLE-Net (adaptive) at $P_{t}=$23\text{\,}\mathrm{d}\mathrm{B}\mathrm{m}$\$ across all scenes.

### v3

# Qiushi Liang, Yeyue Cai, Jianhua Mo, and Meixia Tao

@@ I Introduction
The evolution toward 6G networks features the deep integration of sensing and communication, giving rise to the paradigm of Integrated Sensing and Communication (ISAC) [6]. Leveraging large bandwidths and massive antenna arrays, ISAC enables high-precision localization as a native network service [8]. However, this integration exposes a dichotomy in how wireless systems interact with multipath propagation. For communications, multipath is a valuable resource that enables diversity and multiplexing gains [1]; for monostatic sensing, it is often regarded as interference that causes ambiguities and degrades performance [9, 13].
@@
This challenge has spurred a paradigm shift from multipath mitigation to multipath exploitation, where multipath components are reinterpreted as rich sources of information. In multipath-rich environments, echoes via non-line-of-sight (NLoS) paths provide additional observation perspectives and spatial diversity, improving sensing accuracy, particularly for targets with fluctuating radar cross-sections (RCS) [22, 15]. For extended targets such as vehicles, multipath echoes can even exceed direct reflections and carry more precise motion information [23]. The goal is thus to transform multipath “from enemy into friend” by exploiting its embedded geometric cues for localization tasks [12].
@@
A highly efficient physical-layer mechanism for spatial exploration in wideband ISAC systems is frequency-dependent beamforming, which forms “rainbow beams” through controlled beam squint[21, 14, 4, 5, 25, 24]. By employing true-time-delay (TTD) devices, often as part of a phase-time array (PTA) [20, 16, 2, 11, 18], this technique uniquely maps different frequency subcarriers to distinct spatial locations, enabling a wide angular sector to be scanned within a single OFDM symbol. This “single-scan” operation reduces overhead and latency, catering to real-time applications. The beam squint effect, usually viewed as a communication impairment, is thus harnessed to encode angular information directly in the frequency domain.
@@
Despite this efficiency, interpreting the complex multipath-rich echoes from rainbow beams remains challenging [17]. The superposition of numerous propagation paths forms a unique frequency-domain “fingerprint” for each target, too intricate for conventional estimators that rely on dominant LoS paths. This complexity, however, creates an opportunity for data-driven methods. Deep learning (DL) excels at modeling non-linear mappings from high-dimensional signals to geometric outputs [19, 3]. While prior works apply DL for DoA estimation [3] or near-field localization with fixed rainbow beams [10], jointly leveraging multipath propagation and adaptive beam squint within a unified neural framework remains largely unexplored.
@@
Most existing methods, including DL-based ones, rely on fixed, pre-designed rainbow beams that are non-adaptive and often suboptimal for specific environments [10]. To overcome this limitation, we adopt a co-design paradigm that jointly optimizes the sensing beam and estimation network [7]. By treating the beamforming parameters—phase-shifter (PS) and TTD values—as learnable weights in an end-to-end framework, the system autonomously shapes its rainbow beams and exploits environmental multipath information. Building on this idea, we propose a novel deep learning framework for single-target localization that jointly optimizes the parameters of a TTD-based rainbow beamformer and a subsequent convolutional neural network (CNN) for position estimation. Our key contributions are summarized as follows:
@@
  • Synergistic Co-design for Adaptive Sensing: We propose a fully differentiable, end-to-end model, named MARBLE-Net (Multipath-Aware Rainbow Beam LEarning Network), that jointly optimizes a learnable rainbow beamformer and a CNN-based location prediction network. This co-design enables the system to adapt its physical-layer sensing beam to the specific propagation environment via gradient-based optimization.
  • Structured Training Methodology: We propose a three-stage training strategy that (i) initializes the learnable rainbow beamformer ($\mathcal{M}_{1}$) by maximizing the total received signal power while keeping the prediction network ($\mathcal{M}_{2}$) fixed to increase the uplink SNR, (ii) alternates between updating $\mathcal{M}_{2}$ and adapting $\mathcal{M}_{1}$ to minimize the localization error, thereby steering the beam pattern to exploit geometrically informative multipath components, and (iii) performs end-to-end joint fine-tuning of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$.
@@ II-A Rainbow Beamforming
$\mathbf{w}_{m}]_{n}=\frac{1}{\sqrt{N_{r}}}\exp\left(-\mathrm{j}\left(\phi_{n}+2\pi(f_{m}-f_{0})\tau_{n}\right)\right),$ (1)
where $f_{m}$ is the frequency of the $m$-th subcarrier and $f_{0}$ is the carrier frequency, $\phi_{n}$ and $\tau_{n}$ are the phase shift and time delay applied at the $n$-th antenna. In our framework, the phase shift and time delay values are learnable parameters. They are initialized based on the strategy in [14] to span a predefined angular range but are subsequently optimized during training.
@@
where $\mathbf{h}_{m}\in\mathbb{C}^{N_{r}\times 1}$ is the channel vector, $P_{t}$ is the transmit power of UAV, $x_{m}$ is the transmitted symbol, and $\eta_{m}$ is i.i.d. complex additive white Gaussian noise (AWGN). The collection of received samples, $\mathbf{y}=[y_{1},\dots,y_{M}]^{T}\in\mathbb{C}^{M\times 1}$, forms the observation vector.
@@ II-B High-Fidelity Ray-Tracing Channel Generation
    LoS-only Scene: An unobstructed environment dominated by the direct LoS path, serving as a baseline.
@@
    ‘circle’ Scene: Features a semicircular wall (Fig. 2(a)) that confines reflected paths to a narrow azimuth range, generating uni-directional multipath with low angular spread.
@@
    ‘rounded-L’ Scene: Composed of a quarter-circle arc and a straight wall (Fig. 2(b)). This structure introduces diverse arrival directions, generating multipath with high angular spread.
@@
    ‘L’ scene: Consists of two perpendicular walls representing an urban street corner. This geometry generates a rich multipath profile including higher-order reflections (e.g., double-bounce paths).
For all scenes, the radio material properties are assigned to be concrete for the ground and brick for the walls. The Blender scene is imported into Sionna, and for each of the 100 000100\,000 random UAV positions, Sionna’s path solver traces interactions to compute all valid propagation paths and synthesize the final frequency-domain channel matrix $\mathbf{H}$.
@@ III-A Learnable Rainbow Beamformer ($\mathcal{M}_{1}$)
This module acts as the physical-layer front end. Its trainable parameters consist solely of the PS and TTD vectors, $\bm{\phi},\bm{\tau}\in\mathbb{R}^{N_{r}}$. Given the channel matrix, it computes the received spectrum $\mathbf{y}$ using (1) and (2), and produces the log-magnitude power spectrum $\mathbf{P}_{\text{dB}}\in\mathbb{R}^{M}$ with each element defined as $\mathbf{P}_{\text{dB},m}=10\log_{10}(|y_{m}|^{2}+\epsilon)$.
@@ III-B Prediction Network ($\mathcal{M}_{2}$)
The prediction network is a 1D CNN designed to regress the target’s 2D coordinates from the 1D power spectrum.
@@ III-C Multi-Stage Training Strategy
To address the non-convex loss landscape and ensure stable convergence, we propose a multi-stage strategy transitioning from physical-layer initialization to end-to-end optimization.
  • Stage 1: Power-Based Beamformer Initialization. We optimize the PS and TTD parameters of $\mathcal{M}_{1}$ to maximize the total received signal power $\sum_{m}{\mathbf{P}_{\text{dB,m}}}$, while keeping $\mathcal{M}_{2}$ fixed. This guarantees a high-SNR input, preventing the network from fitting to noise during the initial phase.
  • Stage 2: Alternating Localization Optimization. We iteratively update the prediction network $\mathcal{M}_{2}$ and the beamformer $\mathcal{M}_{1}$ in an alternating manner. Crucially, the beamformer is now updated to minimize the localization error rather than maximizing power, implicitly steering the beam pattern to exploit geometrically informative multipath components.
  • Stage 3: End-to-End Joint Fine-tuning. Finally, all parameters are unfrozen to jointly train the entire MARBLE-Net architecture.
@@ IV-A Simulation Setup
Scene Metric k-NN RaiNet [10] RaiNet MARBLE-Net MARBLE-Net
@@
LoS-only Localization RMSE (m) 15.73 3.72 1.96 1.70 1.52
Angle RMSE () 0.13 1.18 1.26 1.28 2.42
Range RMSE (m) 15.73 3.63 1.89 1.66 1.44
‘circle’ Localization RMSE (m) 16.63 4.12 2.85 2.84 2.30
Angle RMSE () 0.18 1.22 1.14 2.55 1.39
Range RMSE (m) 16.63 4.01 2.76 2.79 2.26
‘rounded-L’ Localization RMSE (m) 13.32 3.76 2.55 2.24 2.15
Angle RMSE () 0.17 1.21 1.12 0.72 0.70
Range RMSE (m) 13.32 3.56 2.27 2.14 2.04
‘L’ Localization RMSE (m) 5.85 1.33 0.86 0.79 0.61
Angle RMSE () 0.14 0.62 0.41 0.25 0.17
Range RMSE (m) 5.84 1.21 0.78 0.74 0.58
@@ IV-B2 RaiNet
As a point of comparison, the baseline RaiNet given in [10] adopted a fixed rainbow beam and only trained the prediction network. This represents a conventional DL-based approach with a non-adaptive sensing beam. The RaiNet architecture consists of a feature extractor and a classifier. The feature extractor contains three 1D convolutional layers, followed by a classifier with three fully connected layers. The Tanh function is used as the primary activation function throughout the network.
@@ IV-C Performance Comparison and Ablation Study
First, it is evident that all deep learning approaches outperform the traditional k-NN baseline. In the ‘circle’ scene, for instance, k-NN yields a large localization RMSE of $16.63\text{\,}\mathrm{m}$, underscoring its inability to handle the complex signal fingerprints created by multipath propagation. Template-matching methods like k-NN fail when the direct relationship between the signal spectrum and geometric location is severely distorted, rendering the pre-compiled codebook ineffective.
@@
To deconstruct the sources of performance gain, we compare four DL-based configurations: RaiNet with a fixed beamformer (‘RaiNet (fixed)’), RaiNet with joint training where the beamformer is adaptive to the scene (‘RaiNet (adaptive)’), MARBLE-Net’s architecture with a fixed beamformer (‘MARBLE-Net (fixed)’), and a scene-adaptive beamformer (‘MARBLE-Net (adaptive)’). This allows us to isolate the effects of network architecture and joint training.
The Benefit of Joint Training: Table II highlights the critical importance of joint end-to-end training between the beamformer and the localization network. Comparing the ‘fixed’ and ‘adaptive’ configurations reveals consistent performance improvements across all scenes. For MARBLE-Net, joint training reduces the localization RMSE by 11%, 19%, 4%, and 23% in the LoS-only, ‘circle’, ‘rounded-L’, and ‘L’ scenes, respectively, while RaiNet achieves even larger relative reductions of 47%, 31%, 32%, and 35%. These results demonstrate that enabling beamformer adaptation with respect to the environment is essential for effectively exploiting multipath information.
The relative gain from joint training differs across architectures. The simpler RaiNet benefits more significantly from beamformer adaptation due to its limited representational capacity and higher sensitivity to the quality of physical-layer observations. In contrast, MARBLE-Net already attains strong performance with a fixed beamformer, resulting in smaller but consistent additional gains from joint optimization. This confirms that network capacity and adaptive beamforming play complementary roles in aligning the sensing strategy with the underlying geometry of multipath propagation.
@@
The Nuanced Impact of Multipath: Table II further reveals a non-monotonic relationship between multipath propagation and localization accuracy. Compared to the LoS-only case, performance degrades in the highly complex ‘circle’ scene, where dense and uni-directional multipath introduces severe geometric ambiguities. For MARBLE-Net, the localization RMSE increases from $1.52\text{\,}\mathrm{m}$ to $2.30\text{\,}\mathrm{m}$, indicating that such confined multipath remains challenging to resolve even with adaptive beamforming.
In contrast, environments with more structured and directionally diverse multipath exhibit improved localization performance. In the ‘rounded-L’ scene, directional NLoS components provide additional spatial diversity, while in the ‘L’ scene, two perpendicular planar walls generate stable and geometrically informative specular reflections. Consequently, the localization RMSE of MARBLE-Net is reduced to $0.61\text{\,}\mathrm{m}$, outperforming all other scenarios, including the LoS-only case. This demonstrates that structured NLoS paths can be effectively exploited as virtual anchors through joint optimization of sensing beam and the location network.
The Benefit of Network Complexity: Table II also shows that increasing network capacity yields clear performance gains. With a fixed beamformer, the more complex MARBLE-Net architecture (‘MARBLE-Net (fixed)’) consistently outperforms ‘RaiNet (fixed)’ across all scenes, for example reducing the localization RMSE from $3.72\text{\,}\mathrm{m}$ to $1.70\text{\,}\mathrm{m}$ in the LoS-only scene. This reflects the stronger capability of a high-capacity network to model the complex signal–location relationships induced by multipath propagation.
Nevertheless, network complexity alone is insufficient to fully exploit multipath-rich environments. The most significant performance improvements are achieved when a high-capacity prediction network is jointly optimized with a learnable, scene-adaptive beamformer, as evidenced by the superiority of ‘MARBLE-Net (adaptive)’.
Figure 5: Localization error CDF of MARBLE-Net at $P_{t}=23$ dBm.
Figure 6: Localization RMSE of MARBLE-Net vs. $P_{t}$.